# Holdout-Loss-Based Data Selection for LLM Finetuning via In-Context Learning

### v1

@@ 5.2 Main Results
When compared to reweighting using baseline scoring methods, our approach achieves consistent gains over one-shot learning across all settings, with win rates consistently above 50%, and often exceeding 60%. Against RHO-Loss, which approximates the same selection criterion in equation 4 but requires training an auxiliary model, our method achieves comparable performance, in some cases surpassing it, while avoiding the cost of training a reference model. These patterns are consistent across both LLaMA and Qwen families and hold for models of different sizes (3B and 8B).
Table 1: Win rates of our method for SFT across models and datasets (higher values indicate better performance), comparing outputs from our method to each baseline.
Alpaca Yahoo_Answers_Topic
Ours against w/o RHO-Loss One-Shot w/o RHO-Loss One-Shot
@@
Table 2: Win rates of our method for DPO across models and datasets (higher values indicate better performance), comparing outputs from our method to each baseline.
@@
Ours against w/o RHO-Loss One-Shot w/o RHO-Loss One-Shot
@@
Table 3: Win rates of our method for simPO across models and datasets (higher values indicate better performance, comparing outputs from our method to each baseline.
@@
Ours against w/o RHO-Loss One-Shot w/o RHO-Loss One-Shot
@@ 5.3 Ablation Studies
We perform ablation studies on LLaMA-3B-Instruct trained on Yahoo_Answers_Topics to examine the effect of key design choices in our method on model performance, with results detailed in Appendix B.6. Key findings are presented below.
@@
To reduce computational cost, we use the top-$k$ holdout examples selected via kNN in embedding space as in-context demonstrations, rather than the full holdout set. Using $k=3$ as the default, smaller or larger values yield no improvement, with win rates below 50% relative to the default: 43.5% for $k=1$, 48.0% for $k=5$, and 46.0% for $k=10$. This indicates that a small number of holdout examples is sufficient to maintain strong alignment while preserving computational efficiency.
@@
We investigate the effect of the total number of score computations $R$ on model performance. In the default setting, scores are computed once at initialization ($R=1$). Increasing $R$ to 3, 5, and 9 improves alignment, with win rates of 50.73%, 52.77%, and 51.6%, respectively. This suggests that additional score updates can further enhance the alignment performance of our method.
@@
We compare percentile-based filtering, which retains examples above a specified percentile, at different thresholds to the default reweighting method. Simple filtering is generally less effective, with win rates below 50% relative to reweighting. A threshold of the 75th percentile yields a higher win rate (48.67%) than the 50th (40.80%) and 90th (40.07%). These results indicate that retaining some lower-scoring examples can be beneficial, but too many degrade performance, so the threshold must be chosen carefully. Adaptive reweighting eliminates this need for manual selection of filtering threshold by automatically adjusting example importance.
@@ 5.4 Analysis
Our method introduces two sources of overhead beyond standard fine-tuning: (i) a one-time precomputation to obtain embeddings, and (ii) periodic score updates. The latter is dominated by the in-context approximation term, which involves additional forward passes and accounts for most of the computational cost. We measure runtime on four NVIDIA A6000 GPUs (48 GB each) for LLaMA-3B-Instruct fine-tuned on Yahoo_Answers_Topics. Precomputation is negligible, on the order of seconds, and the additional time for score computation and reweighting is reported as a percentage relative to the runtime of standard fine-tuning. Our method adds only $\sim$1.5% overhead, compared to roughly 10% for RHO-Loss and 4% for one-shot learning. Full results are provided in Appendix B.7.
@@
We analyze SFT on Yahoo_Answers_Topic to examine the scores assigned to training examples. Using a holdout set sampled from the Sports topic, we compute the average score for examples from each domain (Figure 1). Examples from Sports receive higher scores under both our method and RHO-Loss compared to One-Shot Learning and other topics, demonstrating that our method achieves strong alignment while requiring less computational overhead. Additional analyses in Appendix B.10 examine response patterns produced by models trained with our method, and Appendix B.11 presents instruction-response pairs with high and low ICA scores, offering further insight into our method’s effectiveness.
@@ 6 Conclusion and Discussion
We propose a theoretically grounded, resource-efficient framework for data selection in LLM fine-tuning. Our approach leverages an in-context approximation (ICA) to estimate the holdout loss of the model after including each candidate example in training, without requiring additional finetuning or a reference model. The resulting ICA scores are used to dynamically reweight gradient updates during fine-tuning. Empirical results show that ICA score-based reweighting consistently improves model alignment across SFT, DPO, and SimPO over diverse datasets and backbone architectures, with only marginal computational overhead ($\sim 1.5\%$).
We also note limitations and directions for future work. Our method relies on a high-quality holdout set as a proxy for the test distribution; noisy or unrepresentative holdouts may reduce generalization to unseen data, even if alignment appears strong. Additionally, ICA is inherently off-policy, so applying it to on-policy methods such as PPO would require frequent recomputation of scores for all examples as new data is generated and model parameters continuously evolve, creating a computational bottleneck. Addressing these challenges is a promising direction for future work.
Appendix A Derivation of Holdout Loss Approximation Under a Bayesian Framework
@@
The holdout loss of the model trained with a particular example included in the training set is approximated as follows:
@@
where the first equality applies Bayes’ rule, the second uses a conditional independence assumption, and the last line drops the candidate-independent term.
@@ B.1 Dataset Splits
  • Yahoo_Answers_Topics The holdout and test sets each contain 3,000 examples from the Sports domain. The training set consists of 1,000 examples from each of the remaining domains, together with the holdout examples, totaling 12,000 training examples.
@@
B.2 Evaluation Stability Across Runs
Each model in the main experiments is trained and evaluated only once. To verify evaluation stability, we repeated the GPT-based evaluation multiple times for LLaMA-3-8B-Instruct on Yahoo_Answers_Topics. Table 4 reports the average win rates and standard deviations across these runs. Higher win rates indicate better alignment, and the small standard deviations show that the evaluation results are stable across runs.
Table 4: Win rates of our method against each baseline for LLaMA-3-8B-Instruct on Yahoo_Answers_Topics, averaged over multiple GPT-based evaluations, with corresponding standard deviations reported.
@@ B.2 Evaluation Stability Across Runs
Win Rate (% ↑) 85.10 54.03 66.93
Std. (% ↓) 0.3 0.2 0.5
@@ B.5 Implementation Details of Baselines
We provide additional details on how the scores are computed for each baseline. These scores are used in place of the ICA score within our reweighting framework.
@@
B.6 Ablation Studies
Tables 69 summarize the ablation studies on LLaMA-3B-Instruct trained on Yahoo_Answers_Topics. Win rates indicate the percentage of responses preferred compared to the default setting ($k=3$, $R=1$, using all-mpnet-base-v2 as the embedding model). Each table corresponds to one ablation dimension: Top-$k$ (Table 6); total number of score computations $R$ during training (Table 7); percentile threshold for filtering (Table 8); and embedding model (Table 9). Higher win rates indicate better performance.
Table 6: Effect of Top-$k$ on win rates; higher values indicate better performance.
Top-$k$ 1 3 5
Win rate (% $\uparrow$) 43.50 48.03 46.07
Table 7: Effect of total number of score computations $R$ on win rates; higher values indicate better performance.
@@ B.6 Ablation Studies
Win rate (% $\uparrow$) 50.73 52.77 51.60
Table 8: Effect of percentile threshold for filtering on win rates; higher values indicate better performance. Here, $\geq x$ indicates retaining examples with scores above the $x$-th percentile.
@@
Win rate (% $\uparrow$) 40.80 48.67 40.07
Table 9: Effect of embedding model on win rates; higher values indicate better performance.
Embedding Model Win rate (% $\uparrow$)
@@ B.7 Computational Overhead
B.8 Additional Results Using LoRA
To assess the effectiveness of our method across different parameter updating paradigms, we conduct LoRA training using LLaMA-3-8B-Instruct. Table 11 reports win rates for LoRA with our reweighting method relative to regular LoRA training across various alignment methods and datasets. Each value represents the percentage of responses judged closer to the target by GPT-4o_2024-08-06, with higher values indicating better performance. The results demonstrate that our method consistently improves alignment, even under the LoRA parameter updating setting.
Table 11: Win rates of LoRA trained with our method relative to LoRA trained without our method using LLaMA-3-8B-Instruct. Higher values indicate better performance.
@@ B.8 Additional Results Using LoRA
Win rate (% $\uparrow$) 71.55 68.13 61.23 57.33 56.03 60.13

### v2

@@ 5.2 Main Results
Compared to baseline reweighting methods, our approach consistently outperforms one-shot learning across all settings, with win rates above 50%, and often exceeding 60%. Against RHO-Loss, which approximates the same selection criterion in equation 4 but requires training an auxiliary reference model, our method achieves comparable—and in some cases superior—performance, while avoiding the additional cost of reference model training. These results hold for both SFT and preference-based training paradigms and across LLaMA and Qwen models at different scales (3B and 8B).
For SFT on the Yahoo_Answers_Topics dataset, and using the same data selection scheme, our method achieves comparable performance to LESS and GREATS, with win rates slightly above 50% on both 8B LLaMA and Qwen models, while being more computationally efficient (see Appendix C.3 for full results). This efficiency advantage arises because LESS and GREATS rely on local-linearization–based influence functions, whose first-order assumptions can be violated under batch data selection and require additional computation to control approximation error, whereas our method avoids this overhead by deriving scores from a Bayesian formulation rather than a first-order expansion.
Table 1: Win rates (%) of our method against each baseline when applied to SFT across models and datasets (higher values indicate better performance of our method).
Alpaca Yahoo_Answers_Topics
Model w/o RHO-Loss One-Shot w/o RHO-Loss One-Shot
@@
Table 2: Win rates (%) of our method against each baseline when applied to DPO across models and datasets (higher values indicate better performance of our method).
@@
Model w/o RHO-Loss One-Shot w/o RHO-Loss One-Shot
@@
Table 3: Win rates (%) of our method against each baseline when applied to simPO across models and datasets (higher values indicate better performance of our method).
@@
Model w/o RHO-Loss One-Shot w/o RHO-Loss One-Shot
@@ 5.3 Ablation Studies
We conduct ablation studies under the SFT setting on LLaMA-3B-Instruct trained on the Yahoo_Answers_Topics dataset to examine the effect of key design choices in our method on model performance. Results are detailed in Appendix B.6, with key findings presented below.
@@
To reduce computational cost, we use the top-$k$ holdout examples selected via kNN in embedding space as in-context demonstrations, rather than the full holdout set. Using $k=3$ as the default, smaller or larger values yield no improvement, with win rates below 50% relative to the default: 43.5% for $k=1$, 48.0% for $k=5$, and 46.0% for $k=10$. The performance drop for $k>3$ may be due to additional holdout examples being less relevant. This indicates that a small number of holdout examples is sufficient to maintain strong alignment while preserving computational efficiency.
@@
We investigate the effect of the total number of score computations $R$ on model performance. In the default setting, scores are computed once at initialization ($R=1$). Increasing $R$ to 3, 5, and 9 yields win rates of 50.73%, 52.77%, and 51.6%, respectively. These results suggest that increasing $R$ can further improve alignment, while very frequent updates (e.g., $R=9$) appear unnecessary, as the dynamics of ICA weights (see Section 5.5) show that the relative importance of training examples stabilize in later stages of training.
@@
As an alternative to our default reweighting approach, we consider percentile-based filtering using ICA scores, retaining only examples above a specific percentile. Simple filtering is generally less effective, with win rates below 50% relative to reweighting. A threshold of the 75th percentile yields a higher win rate (48.67%) than the 50th (40.80%) and 90th (40.07%), indicating that retaining some lower-scoring examples can be beneficial, while too many degrade performance, so the threshold must be chosen carefully. In contrast, adaptive reweighting eliminates this need for manually selecting a filtering threshold by automatically adjusting example importance.
@@
5.4 Additional Results
We present additional results to complement the main evaluations, using the SFT setting as a representative example. Specifically, we report absolute performance metrics that do not rely on a model-based judge and further evaluate the effectiveness of our method on a reasoning-intensive benchmark. We highlight the main findings below, with detailed results deferred to Appendix C.2 and C.4.
Consistency across metrics
To complement GPT-judged win rates, we report absolute performance as measured by perplexity (PPL) and BERTScore (Zhang et al., 2019) for SFT models trained on Yahoo_Answers_Topics using our reweighting method, compared against each baseline. As shown in Appendix C.2, the performance trends under these absolute metrics are consistent with those observed under GPT-judged win rates.
Mitigating the effect of noise on GSM8k
To evaluate our method’s ability to select valuable training examples from datasets that may contain suboptimal instances, we conduct SFT on both the original training split and a variant with intentionally corrupted CoT annotations. Details on how the data were corrupted and how the holdout set was constructed are provided in Appendix B.1. Pass@1 accuracy under zero-shot CoT prompting is reported in Table 15 in Appendix C.4, with key findings summarized below.
SFT trained on the original high-quality split achieves the best performance. Using our reweighting method on the corrupted split yields the second-best results, with only a small performance gap. These results show that reweighting noisy training data with our method can achieve performance comparable to training on the original high-quality dataset, highlighting its potential to mitigate the impact of substantial noise, even on complex reasoning tasks.
@@ 5.5 Analysis
Our method introduces two sources of overhead beyond standard fine-tuning: (i) a one-time precomputation to obtain embeddings, and (ii) periodic score updates. The latter is dominated by the ICA term, which involves additional forward passes and accounts for most of the computational cost. We measure runtime on four NVIDIA A6000 GPUs (48 GB each) for LLaMA-3B-Instruct fine-tuned on Yahoo_Answers_Topics. Precomputation is negligible, on the order of seconds, and the additional time for score computation and reweighting is reported as a percentage of standard fine-tuning runtime. Our method adds only $\sim$1.5% overhead, compared to roughly 10% for RHO-Loss and 4% for one-shot learning. Full results are provided in Appendix B.7.
@@
We analyze SFT on Yahoo_Answers_Topics to examine the scores assigned to training examples. Using a holdout set sampled from the Sports topic, we compute the average score for examples from each domain (Figure 1). Examples from Sports receive higher scores under both our method and RHO-Loss compared to One-Shot Learning and other topics, demonstrating that our method achieves strong alignment while requiring less computational overhead.
Dynamics of ICA weights
We observe that ICA weights adjust dynamically during the training process, with most changes occurring early and stabilizing as training progresses. Using SFT results on the Yahoo_Answers_Topics dataset with $R=5$ as an example, correlations between weights from the first update and subsequent updates are 0.89, 0.75, 0.69, and 0.71. Since the weights stabilize later, very frequent score updates are unnecessary. Additional analyses in Appendix D.2 examine response patterns produced by models trained with our method, and Appendix D.3 presents instruction-response pairs with high and low ICA scores, offering further insight into our method’s effectiveness.
@@ 6 Conclusion and Discussion
We propose a principled, resource-efficient framework for data selection in LLM fine-tuning. Our approach leverages an in-context approximation (ICA) to estimate the holdout loss of the model after including each candidate example in training, without requiring additional finetuning or a reference model. The resulting ICA scores are used to dynamically reweight gradient updates during fine-tuning. Empirical results show that ICA score-based reweighting consistently improves model alignment across SFT, DPO, and SimPO over diverse datasets and backbone architectures, with only marginal computational overhead ($\sim 1.5\%$).
We also note limitations and directions for future work. We focus on settings where a high-quality holdout set is available to estimate the importance of training examples. If the holdout set is noisy or unrepresentative, generalization to unseen data may be impaired, even if alignment appears strong. Additionally, our experiments use off-policy training paradigms. Directly applying ICA to on-policy methods, such as PPO, would require frequent recomputation of scores for all newly generated data, creating a substantial computational bottleneck. Addressing these challenges is a promising direction for future work.
Appendix A Derivation of equation 3 Under a Bayesian Framework
@@
Under a Bayesian framework, the holdout loss of a model trained with a particular example can be expressed as
@@
where the first equality applies Bayes’ rule, the second uses a conditional independence assumption, i.e., $p(\mathbf{y}_{i}\mid\mathbf{x}_{i},\mathbf{x}_{j},\mathcal{D}_{t})=p(\mathbf{y}_{i}\mid\mathbf{x}_{i},\mathcal{D}_{t})$, for any $i\neq j$, and the last line drops the candidate-independent term.
@@ B.1 Dataset Splits
  • Yahoo_Answers_Topics The holdout and test sets each contain 3,000 examples from the Sports domain. The training set consists of 10,000 examples from each of the remaining domains, together with the holdout examples, totaling 100,300 training examples.
@@
  • GSM8K Cobbe et al. (2021) This dataset contains 8.5K high-quality grade school math word problems requiring multi-step reasoning. From the training split, we set aside 1,000 examples as a high-quality holdout. The remaining training examples are used as the model training set, of which a randomly selected 40% are intentionally corrupted (e.g., via CoT dropout, shuffled CoT, or CoT re-generation using a weaker model). The original test split is used for evaluation.
In main experiments, the holdout set is included in the training data to ensure fair comparison with baselines that do not use reweighting. We additionally conduct experiments where the holdout set is used only for scoring in Appendix C.5.
B.2 Training and Evaluation Variability
Each model in the main experiments is trained and evaluated only once. To verify evaluation stability, we repeated the GPT-based evaluation three times for LLaMA-3-8B-Instruct on Yahoo_Answers_Topics. Table 4 reports the average win rates of our method against each baseline, along with standard deviations across these runs. Higher win rates indicate better alignment of our method, and the small standard deviations show that the evaluation results are stable across runs.
To assess training variability, we performed three independent SFT runs with different random seeds for each of the LaMA-3-8B-Instruct and Qwen-3-8B models on Yahoo_Answers_Topics. Table 5 reports the average win rates of our method against each baseline, with standard deviations provided for reference. The mean performance across these training runs aligns closely with the single-run results reported in the main experiments.
Table 4: Win rates of our method against each baseline for LLaMA-3-8B-Instruct on Yahoo_Answers_Topics, averaged over three GPT-based evaluations from a single training run, with corresponding standard deviations reported.
@@ B.2 Training and Evaluation Variability
Win Rate (%) 85.10 $\pm$ 0.30 54.03 $\pm$ 0.20 66.93 $\pm$ 0.50
Table 5: Win rates (%) of our method against each baseline on Yahoo_Answers_Topics, averaged over three independent runs with different random seeds, reported for both LLaMA-3-8B-Instruct and Qwen-3-8B. Standard deviations are also provided for reference.
Model w/o RHO-Loss One-shot
LLaMA 8B 82.86 $\pm$ 1.96 49.92 $\pm$ 0.39 62.59 $\pm$ 0.57
Qwen 8B 81.67 $\pm$ 1.09 54.23 $\pm$ 0.18 63.81 $\pm$ 0.60
@@ B.5 Implementation Details of Baselines
We provide additional details on how scores are computed for each reweighting baseline. These scores are used in place of the ICA score within our reweighting framework. For the subset selection methods LESS Xia et al. (2024) and GREATS Wang et al. (2024), we reimplement their publicly available code and follow their original data selection procedures: selecting the top 5% before training for LESS, and 50% of each batch for gradient updates for GREATS.
@@
B.6 Results of Ablation Studies
Tables 710 summarize the ablation studies on LLaMA-3B-Instruct trained on Yahoo_Answers_Topics. Win rates indicate the percentage of responses preferred over the default setting (i.e., $k=3$, $R=1$, using all-mpnet-base-v2 as the embedding model). Each table corresponds to varying one design dimension: Top-$k$ (Table 7); total number of score computations $R$ during training (Table 8); percentile threshold for filtering (Table 9); and embedding model (Table 10). Higher win rates indicate better performance of the corresponding design choice; 50% corresponds to parity with the default.
Table 7: Effect of Top-$k$ on win rates; higher values indicate better performance for the corresponding $k$.
Top-$k$ 1 5 10
Win rate (%) 43.50 48.03 46.07
Table 8: Effect of total number of score computations $R$ on win rates; higher values indicate better performance for the corresponding $R$.
@@ B.6 Results of Ablation Studies
Win rate (% ) 50.73 52.77 51.60
Table 9: Effect of percentile threshold for filtering on win rates; higher values indicate better performance for the corresponding percentile. Here, $\geq x$ indicates retaining examples with scores above the $x$-th percentile.
@@
Win rate (%) 40.80 48.67 40.07
Table 10: Win rates under a stronger embedding model against the default; higher values indicate better performance for the stronger model.
Embedding Model Win rate (%)
@@ B.7 Computational Overhead
Appendix C Additional Experiments
C.1 Additional Results Using LoRA
To assess the effectiveness of our method across different parameter updating paradigms, we conduct LoRA training using LLaMA-3-8B-Instruct. Table 12 reports win rates for LoRA with our reweighting method relative to regular LoRA training across various alignment methods and datasets. Each value represents the percentage of responses judged closer to the target by GPT-4o (2024-08-06), with higher values indicating better performance when our method is applied. The results demonstrate that our method consistently improves alignment, including under the LoRA parameter updating setting.
Table 12: Win rates of LoRA trained with our method against a LoRA baseline trained without our method using LLaMA-3-8B-Instruct. Higher values indicate better performance when our method is applied.
@@ C.1 Additional Results Using LoRA
Win rate (%) 71.55 68.13 61.23 57.33 56.03 60.13
C.2 Absolute Metrics on Yahoo_Answers_Topics
Table 13 reports perplexity (PPL) and BERTScore for SFT models trained on Yahoo_Answers_Topics under different reweighting methods. Lower PPL and higher BERTScore indicate better model performance.
Table 13: Absolute-metric performance of SFT-trained models on Yahoo_Answers_Topics under different reweighting strategies (including no reweighting), evaluated using perplexity (PPL) and BERTScore. Lower PPL and higher BERTScore indicate better model performance. Best and second-best results are highlighted in bold and underlined, respectively.
LLaMA 8B Qwen 8B
Method PPL ($\downarrow$) BERTScore ($\uparrow$) PPL ($\downarrow$) BERTScore ($\uparrow$)
Ours 11.23 0.84 10.30 0.87
w/o 16.71 0.80 13.04 0.82
RHO-Loss 11.49 0.84 10.00 0.88
One-shot 14.92 0.82 12.02 0.82
C.3 Comparison with LESS and GREATS
Table 14 reports win rates of our method against LESS and GREATS under the same data selection paradigm for SFT on Yahoo_Answers_Topics, together with PPL and BERTScore for each method. Comparison results on GSM8k are reported in the following section, using Pass@1 accuracy. Runtime of computing scores for all training examples using each data selection method is provided in Table 16.
Table 14: Comparison of our method with additional baselines under the SFT setting on Yahoo_Answers_Topics. Win rates are compared within the same selection paradigm; higher values indicate stronger performance of our method relative to the corresponding baseline. PPL and BERTScore are also reported. with the best values indicated in bold.
LLaMA-3-8B Qwen-3-8B
Method Win rate (Ours vs Baseline) PPL ($\downarrow$) BERTScore ($\uparrow$) Win rate (%) (Ours vs Baseline) PPL ($\downarrow$) BERTScore ($\uparrow$)
Online selection
GREATS (50% per batch) 52.03 22.07 0.78 50.91 19.70 0.80
Ours (50% per batch) 22.01 0.79 19.74 0.80
Pre-training selection
LESS (Top 5%) 51.40 20.55 0.80 50.90 17.24 0.77
Ours (Top 5%) 18.91 0.80 17.03 0.78
C.4 Results on GSM8k
Table 15 reports Pass@1 accuracy for SFT models trained on the original and corrupted splits of GSM8k using our reweighting method and baseline approaches.
Table 15: Pass@1 accuracy (%) on GSM8k under zero-shot CoT prompting for LLaMA-3-8B and Qwen-3-8B, trained with SFT using different data selection strategies. Best and second-best results are highlighted in bold and underlined, respectively.
Method LLaMA-3-8B Qwen-3-8B
Without reweighting
Pretrained 75.28 80.52
SFT (original) 81.91 83.27
SFT (corrupted) 52.75 62.40
Reweighting-based methods
RHO-Loss 79.50 81.16
One-shot 72.42 76.19
Ours (reweighting) 79.95 83.04
Pre-training selection
LESS (Top 5%) 77.63 80.81
Ours (Top 5%) 78.92 80.81
Online selection
GREATS (50% per batch) 80.03 80.50
Ours (50% per batch) 80.18 80.49
Table 16: Score computation time (in seconds) on LLaMA-3-8B for our method, LESS, and GREATS when training on Yahoo_Answers_Topics with SFT. Times for LESS and GREATS are based on our reimplementation. For LESS, the two numbers correspond to influence score computation and periodic model retraining to address the circular dependence between data selection and model updates.
Dataset Ours LESS GREATS
Yahoo_Answers_Topics 808 1,394 + 1,800 1,004
GSM8K 117 310 + 900 704
Table 17: Win rates of our method against each baseline reweighting method for models trained with SFT on Yahoo_Answers_Topics, using $\mathcal{D}_{\mathrm{ho}}$ as a strict holdout for scoring (not included in training). Higher win rates indicate better performance of our method. PPL and BERTScore are also reported, with the best and second-best values indicated in bold and underlined, respectively.
LLaMA-3-8B Qwen-3-8B
Method Win rate (Ours vs Baseline) PPL ($\downarrow$) BERTScore ($\uparrow$) Win rate (%) (Ours vs Baseline) PPL ($\downarrow$) BERTScore ($\uparrow$)
w/o 89.27 26.77 0.72 87.49 22.44 0.80
RHO-Loss 47.29 12.08 0.81 49.21 12.94 0.85
One-shot 71.33 19.45 0.79 67.02 18.73 0.82
Ours 14.05 0.81 13.29 0.86
Table 18: Comparison of our method against the baseline model trained only on the holdout set $\mathcal{D}_{\mathrm{ho}}$ for SFT on Yahoo_Answers_Topics. Win rates indicate the percentage of examples where our method is preferred over the baseline; higher values indicate better performance of our method. PPL and BERTScore are also reported. with best values indicated in bold.
LLaMA-3-8B Qwen-3-8B
Method Win rate (Ours vs Baseline) PPL ($\downarrow$) BERTScore ($\uparrow$) Win rate (%) (Ours vs Baseline) PPL ($\downarrow$) BERTScore ($\uparrow$)
Trained only on $\mathcal{D}_{\mathrm{ho}}$ 58.13 12.40 0.82 56.80 11.92 0.82
Ours 11.23 0.84 10.30 0.87
C.5 Excluding Holdout Examples from Training
In our main experiments, the holdout set is included in training to ensure a fair comparison with baselines that do not use reweighting, which otherwise have no access to holdout examples. To examine the potential impact of including holdout examples on performance, we reran SFT on the Yahoo_Answers_Topic dataset, keeping the holdout set strictly for scoring. Results are presented in Table 17. Even without training on the holdout set, our method outperforms the standard baseline and one-shot learning, although it performs slightly below RHO-Loss, which trains a reference model on the holdout set. Nevertheless, our method remains comparable to RHO-Loss while using holdout examples only as in-context demonstrations for scoring.
C.6 Performance When Using the Holdout Set Alone
To show that test performance relies on the broader training data and the holdout set alone does not suffice, we train the model solely on the holdout set (see results in Table 18). As expected, its performance is lower than our method trained on the full training set with ICA reweighting, highlighting the benefit of combining broader training data with data selection guided by the holdout set.
Appendix D Additional Analysis